ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 35A30, 35K25 Shimomura, Katsunori, Ibaraki University, Japan On the transformations which preserve polytemperatures A Polytemperature of degree m is a solution of the mth-iterate of the heat operator. We treat the transformations, pairs of a mapping and a positive multiplier, which preserve polytemperatures of degree m. A necessary and sufficient conditions will be given. Also a close connection between such transformations and caloric morphisms will be presented.